OpenStudy (anonymous):

What is an equation for the line that passes through points (–1, –4) and (1, 4)? Write the equation in slope-intercept form. y = 4x y = –4x y = –1x – 4 y = –4x + 1

OpenStudy (hba):

Do you like one formula or two formulas to work this out ?

OpenStudy (anonymous):

two please

OpenStudy (anonymous):

First find the slope of these two points, so (4 - (-4))/(1 - (-1)).. The slope is 8/2 or 4. Then you can just plug in a point to find b, into the equation y = 4x + b, so you can do (4) = 4(1) + b. We get b = 0 so the answer is y = 4x.

OpenStudy (hba):

Well then lets start using two formulas, First of all we will find the slope, We know that, \[\huge\ m=\frac{ y_2-y_1 }{ x_2-x_1 }\] Here, (x1,y1)= (–1, –4) and (x2,y2)= (1, 4)

OpenStudy (anonymous):

Thank you both!